# Each of those eight coalition logics is also determined by six other kinds of models

### v2

@@ 5.1 Clear grand-coalition-first action models
    for all $\mathrm{C}\subseteq\mathtt{AG}$, $s\in\mathtt{ST}$, and $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$,
    $\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})=\begin{cases}\mathtt{suc}(s)&\text{if $\mathrm{C}=\emptyset$}\\ \bigcap\{\mathsf{out}_{a}(s,\sigma_{{a}})\mid a\in\mathrm{C},\sigma_{{a}}\in\mathtt{JA}_{a},\text{and }\sigma_{{a}}\subseteq\sigma_{{\mathrm{C}}}\}&\text{if $\mathrm{C}\neq\emptyset$}\end{cases}$
  3. (3)
    for every $\mathrm{C}\subseteq\mathtt{AG}$, for all $s\in\mathtt{ST}$,
    $\mathtt{av}_{\mathrm{C}}(s)=\{\sigma_{{\mathrm{C}}}\in\mathtt{JA}_{\mathrm{C}}\mid\mathsf{out}(s,\sigma_{{\mathrm{C}}})\neq\emptyset\}.$
By Fact 2, (1) holds. Clearly, (3) holds. It remains to show (2). Let $\mathrm{C}\subseteq\mathtt{AG}$, $s\in\mathtt{ST}$, and $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$.
Assume $\mathrm{C}=\emptyset$. Then, $\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})=\mathsf{out}_{\emptyset}(s,\emptyset)=\bigcup\{\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})\mid\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}\text{ and }\emptyset\subseteq\sigma_{\mathtt{AG}}\}=\bigcup\{\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})\mid\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}\}=\mathtt{suc}(s)$;
Assume $\mathrm{C}\neq\emptyset$. Let $\mathrm{C}=\{a_{1},\dots,a_{n}\}$ and $\sigma_{\mathrm{C}}=\sigma_{a_{1}}\cup\dots\cup\sigma_{a_{n}}$. By Fact 3, $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\subseteq\mathsf{out}_{a_{1}}(s,\sigma_{a_{1}})\cap\dots\cap\mathsf{out}_{a_{n}}(s,\sigma_{a_{n}})$. Let $t\in\mathsf{out}_{a_{1}}(s,\sigma_{a_{1}})\cap\dots\cap\mathsf{out}_{a_{n}}(s,\sigma_{a_{n}})$. Then, there is $\sigma^{1}_{\mathtt{AG}}$ such that $\sigma_{a_{1}}\subseteq\sigma^{1}_{\mathtt{AG}}$ and $t\in\mathsf{out}_{\mathtt{AG}}(s,\sigma^{1}_{\mathtt{AG}})$, …, there is $\sigma^{n}_{\mathtt{AG}}$ such that $\sigma_{a_{n}}\subseteq\sigma^{n}_{\mathtt{AG}}$ and $t\in\mathsf{out}_{\mathtt{AG}}(s,\sigma^{n}_{\mathtt{AG}})$. As $\mathtt{G}\textbf{-}\mathtt{AM}$ is clear, $\sigma^{1}_{\mathtt{AG}}=\dots=\sigma^{n}_{\mathtt{AG}}$. Then $\sigma_{\mathrm{C}}\subseteq\sigma^{1}_{\mathtt{AG}}$. Then $t\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})$.
It is easy to find a single-coalition-first action model that is not clear.
@@
Goranko and Jamroga [4] and Ågotnes and Alechina [12] discussed injective concurrent game models, which are clear grand-coalition-first action $\mathtt{SID}$-models.
As observed by Ågotnes and Alechina [12], the assumption behind injective concurrent game models, different available action profiles have different outcomes, is very common in game theory.
@@ 5.2 Clear single-coalition-first neighborhood models
Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model.
We say $\mathtt{S}\textbf{-}\mathtt{NM}$ is clear if for all $a\in\mathtt{AG}$ and $s\in\mathtt{ST}$, $\mathtt{nei}_{a}(s)$ is a partition of $\mathtt{suc}(s)$.
@@
Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model, and $\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the class of neighborhood functions of $\mathtt{S}\textbf{-}\mathtt{NM}$.
@@
    for all $a\in\mathtt{AG}$ and $s\in\mathtt{ST}$, $\mathtt{nei}_{a}(s)$ is a partition of $\mathtt{suc}(s)$;
@@
    for all $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, $\mathtt{nei}_{\mathrm{C}}(s)$ is a partition of $\mathtt{suc}(s)$.
@@
Assume (1). Let $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$. By Fact 4, $\mathtt{nei}_{\mathrm{C}}(s)$ is a cover of $\mathtt{suc}(s)$. Assume $\mathtt{nei}_{\mathrm{C}}(s)$ is not a partition of $\mathtt{suc}(s)$. Then, there is $\mathrm{Y}_{1},\mathrm{Y}_{2}\in\mathtt{nei}_{\mathrm{C}}(s)$ such that $\mathrm{Y}_{1}\neq\mathrm{Y}_{2}$ and $\mathrm{Y}_{1}\cap\mathrm{Y}_{2}\neq\emptyset$. Without loss of any generality, we can assume there is $w$ such that $w\in\mathrm{Y}_{1}$ but $w\notin\mathrm{Y}_{2}$. Let $u\in\mathrm{Y}_{1}\cap\mathrm{Y}_{2}$. It is easy to see $\mathtt{suc}(s)\neq\emptyset$. Then, $\mathtt{nei}_{\emptyset}(s)=\{\mathtt{suc}(s)\}$, which is a singleton. Note $\mathtt{nei}_{\mathrm{C}}(s)$ is not a singleton. Then, $\mathrm{C}\neq\emptyset$. Let $\mathrm{C}=\{a_{1},\dots,a_{n}\}$. Then, there is $\mathrm{X}_{a_{1}}\in\mathtt{nei}_{a_{1}}(s),\dots,\mathrm{X}_{a_{n}}\in\mathtt{nei}_{a_{n}}(s)$ such that $\mathrm{Y}_{1}=\mathrm{X}_{a_{1}}\cap\dots\cap\mathrm{X}_{a_{n}}$, and there is $\mathrm{X}^{\prime}_{a_{1}}\in\mathtt{nei}_{a_{1}}(s),\dots,\mathrm{X}^{\prime}_{a_{n}}\in\mathtt{nei}_{a_{n}}(s)$ such that $\mathrm{Y}_{2}=\mathrm{X}^{\prime}_{a_{1}}\cap\dots\cap\mathrm{X}^{\prime}_{a_{n}}$. Note $w\notin\mathrm{Y}_{2}$. Then $w\notin\mathrm{X}^{\prime}_{a_{i}}$ for some $a_{i}$. Note $w\in\mathrm{X}_{a_{i}}$. Then $\mathrm{X}_{a_{i}}\neq\mathrm{X}^{\prime}_{a_{i}}$. Note $u\in\mathrm{X}_{a_{i}}$ and $u\in\mathrm{X}^{\prime}_{a_{i}}$. Then $\mathrm{X}_{a_{i}}\cap\mathrm{X}^{\prime}_{a_{i}}\neq\emptyset$. We have a contradiction.
@@
By Theorem 9, clear grand-coalition-first action models can be equivalently defined in three different ways, which respectively concern all agents, all coalitions, and the grand coalition. It might be expected that the following condition, which concerns the grand coalition, is equivalent to the two conditions given in Theorem 11: for all $s\in\mathtt{ST}$, $\mathtt{nei}_{\mathtt{AG}}(s)$ is a partition of $\mathtt{suc}(s)$. Actually, this is not the case. What follows is a counterexample.
Assume $\mathtt{AG}=\{a,b\}$. Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model such that:
@@
Note $\mathtt{nei}_{a}(s_{0})$ is not a partition of $\mathtt{suc}(s_{0})$. However, it can be checked that $\mathtt{nei}_{\mathtt{AG}}(s_{0})=\{\{s_{1}\},\{s_{2}\},\{s_{3}\}\}$, which is a partition of $\mathtt{suc}(s_{0})$.
@@
Alurz, Henzinger, and Kupferman [2] proposed models for $\mathsf{ATL}$ based on so-called lock-step synchronous alternating transition systems. It can be shown that they are clear single-coalition-first neighborhood $\mathtt{SID}$-models.
@@
Let $\mathtt{G}\textbf{-}\mathtt{AM}$ be a clear grand-coalition-first action model. By Theorem 10, $\mathtt{G}\textbf{-}\mathtt{AM}$ is a single-coalition-first action model. It is easy to show that $\mathtt{G}\textbf{-}\mathtt{AM}$ is $z$-representable by a clear single-coalition-first neighborhood model. ∎
@@
Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a clear single-coalition-first neighborhood model. Define a single-coalition-first action model $\mathtt{S}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathtt{suc},\{\mathsf{out}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ as in the proof for Theorem 6. As shown in that proof, $\mathtt{S}\textbf{-}\mathtt{NM}$ $z$-represents $\mathtt{S}\textbf{-}\mathtt{AM}$. It is easy to check that $\mathtt{S}\textbf{-}\mathtt{AM}$ is clear. By Theorem 4, it is a grand-coalition-first action model.
@@ 6.1 Tree-like grand-coalition-first action models
Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a grand-coalition-first action model, $\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of availability functions of $\mathtt{G}\textbf{-}\mathtt{AM}$, and $\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of outcome functions of $\mathtt{G}\textbf{-}\mathtt{AM}$.
@@
Specially, for every $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, $s$ is called a $\mathrm{C}$-history from $s$ to $s$.
@@
Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a grand-coalition-first action model.
We say $\mathtt{G}\textbf{-}\mathtt{AM}$ is a tree-like model if there is $r\in\mathtt{ST}$, called a root, such that for every $s\in\mathtt{ST}$, there is a unique $\mathtt{AG}$-history from $r$ to $s$.
Note every tree-like model has a unique root.
@@
Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a grand-coalition-first action model, $\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of availability functions of $\mathtt{G}\textbf{-}\mathtt{AM}$, and $\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of outcome functions of $\mathtt{G}\textbf{-}\mathtt{AM}$.
@@
    there is $r\in\mathtt{ST}$ such that for every $a\in\mathtt{AG}$ and $s\in\mathtt{ST}$, there is an unique $a$-history from $r$ to $s$;
@@
    there is $r\in\mathtt{ST}$ such that for every $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, there is an unique $\mathrm{C}$-history from $r$ to $s$;
@@
    there is $r\in\mathtt{ST}$ such that for every $s\in\mathtt{ST}$, there is an unique $\mathtt{AG}$-history from $r$ to $s$.
@@
Assume (1). Let $r\in\mathtt{ST}$ such that for every $a\in\mathtt{AG}$ and $s\in\mathtt{ST}$, there is an unique $a$-history from $r$ to $s$. Let $s\in\mathtt{ST}$. We want to show that there is an unique $\mathtt{AG}$-history from $r$ to $s$. Let $a\in\mathtt{AG}$.
@@
Let $\theta_{a}$ be the unique $a$-history from $r$ to $s$. Note it is impossible $\theta_{a}=x$ for some $x\in\mathtt{ST}$. Assume $\theta_{a}=(s_{0},\sigma_{a}^{1},s_{1},\dots,\sigma_{a}^{n},s_{n})$.
@@
Assume there is a different $\mathtt{AG}$-history $\theta^{\prime}_{\mathtt{AG}}$ from $r$ to $s$. Note it is impossible $\theta^{\prime}_{\mathtt{AG}}=x$ for some $x\in\mathtt{ST}$. Let $\theta^{\prime}_{\mathtt{AG}}=(t_{0},\lambda^{1}_{\mathtt{AG}},t_{1},\dots,\lambda_{\mathtt{AG}}^{m},t_{m})$.
@@
Assume (3). Let $r\in\mathtt{ST}$ such that for every $s\in\mathtt{ST}$, there is an unique $\mathtt{AG}$-history from $r$ to $s$. Let $s\in\mathtt{ST}$ and $\mathrm{C}\subseteq\mathtt{AG}$. We want to show that there is an unique $\mathrm{C}$-history from $r$ to $s$.
@@
Assume there is a $\mathrm{C}$-history $\theta_{\mathrm{C}}$ from $r$ to $r$ which is different from $r$. Note it is impossible $\theta_{\mathrm{C}}=x$ for some $x\in\mathtt{ST}$. Assume $\theta_{\mathrm{C}}=(s_{0},\sigma_{\mathrm{C}}^{1},s_{1},\dots,\sigma_{\mathrm{C}}^{n},s_{n})$.

### v3

@@ 5.1 Clear grand-coalition-first action models
    for all nonempty $\mathrm{C}\subseteq\mathtt{AG}$, $s\in\mathtt{ST}$, and $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$,
    $\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})=\bigcap\{\mathsf{out}_{a}(s,\sigma_{{a}})\mid a\in\mathrm{C},\sigma_{{a}}\in\mathtt{JA}_{a},\text{and }\sigma_{{a}}\subseteq\sigma_{{\mathrm{C}}}\}$
By Fact 2, (1) holds. It remains to show (2). Let $\mathrm{C}\subseteq\mathtt{AG}$, $s\in\mathtt{ST}$, and $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$. Assume $\mathrm{C}\neq\emptyset$
Let $\mathrm{C}=\{a_{1},\dots,a_{n}\}$ and $\sigma_{\mathrm{C}}=\sigma_{a_{1}}\cup\dots\cup\sigma_{a_{n}}$. By Fact 3, $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\subseteq\mathsf{out}_{a_{1}}(s,\sigma_{a_{1}})\cap\dots\cap\mathsf{out}_{a_{n}}(s,\sigma_{a_{n}})$. Let $t\in\mathsf{out}_{a_{1}}(s,\sigma_{a_{1}})\cap\dots\cap\mathsf{out}_{a_{n}}(s,\sigma_{a_{n}})$. Then, there is $\sigma^{1}_{\mathtt{AG}}$ such that $\sigma_{a_{1}}\subseteq\sigma^{1}_{\mathtt{AG}}$ and $t\in\mathsf{out}_{\mathtt{AG}}(s,\sigma^{1}_{\mathtt{AG}})$, …, there is $\sigma^{n}_{\mathtt{AG}}$ such that $\sigma_{a_{n}}\subseteq\sigma^{n}_{\mathtt{AG}}$ and $t\in\mathsf{out}_{\mathtt{AG}}(s,\sigma^{n}_{\mathtt{AG}})$. As $\mathtt{G}\textbf{-}\mathtt{AM}$ is clear, $\sigma^{1}_{\mathtt{AG}}=\dots=\sigma^{n}_{\mathtt{AG}}$. Then $\sigma_{\mathrm{C}}\subseteq\sigma^{1}_{\mathtt{AG}}$. Then $t\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})$.
It is easy to see that the action model given in Example 5 is a single-coalition-first action model that is not clear.
@@
Goranko and Jamroga [goranko_comparing_2004] and Ågotnes and Alechina [agotnes_embedding_2015] discussed injective concurrent game models, which are clear grand-coalition-first action $\mathtt{SID}$-models. As observed by Ågotnes and Alechina [agotnes_embedding_2015], the assumption behind injective concurrent game models, different available action profiles have different outcomes, is very common in game theory.
The reason that Goranko and Jamroga [goranko_comparing_2004] and Ågotnes and Alechina [agotnes_embedding_2015] used “injective” is as follows: in injective concurrent game models, the outcome function $\mathsf{out}_{\mathtt{AG}}$ for $\mathtt{AG}$ is an injective function with the set of possible states as its range. However, in grand-coalition-first models, the outcome function $\mathsf{out}_{\mathtt{AG}}$ for $\mathtt{AG}$ takes the power set of the set of possible states as its range. A grand-coalition-first action model where the outcome function $\mathsf{out}_{\mathtt{AG}}$ for $\mathtt{AG}$ is injective does not imply that it is clear. This is why we do not use “injective” for clear grand-coalition-first action models.
@@ 5.2 Clear single-coalition-first neighborhood models
Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model.
We say that $\mathtt{S}\textbf{-}\mathtt{NM}$ is clear if for all $a\in\mathtt{AG}$ and $s\in\mathtt{ST}$, $\mathtt{nei}_{a}(s)$ is a partition of $\bigcup\mathtt{nei}_{\emptyset}(s)$.
@@
Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model.
@@
    for all $a\in\mathtt{AG}$ and $s\in\mathtt{ST}$, $\mathtt{nei}_{a}(s)$ is a partition of $\bigcup\mathtt{nei}_{\emptyset}(s)$;
@@
    for all $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, $\mathtt{nei}_{\mathrm{C}}(s)$ is a partition of $\bigcup\mathtt{nei}_{\emptyset}(s)$.
@@
Assume (1). Let $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$. By Fact 4, $\mathtt{nei}_{\mathrm{C}}(s)$ is a cover of $\bigcup\mathtt{nei}_{\emptyset}(s)$. Assume $\mathtt{nei}_{\mathrm{C}}(s)$ is not a partition of $\bigcup\mathtt{nei}_{\emptyset}(s)$. Then, there is $\mathrm{Y}_{1},\mathrm{Y}_{2}\in\mathtt{nei}_{\mathrm{C}}(s)$ such that $\mathrm{Y}_{1}\neq\mathrm{Y}_{2}$ and $\mathrm{Y}_{1}\cap\mathrm{Y}_{2}\neq\emptyset$. Without loss of any generality, we can assume there is $w$ such that $w\in\mathrm{Y}_{1}$ but $w\notin\mathrm{Y}_{2}$. Let $u\in\mathrm{Y}_{1}\cap\mathrm{Y}_{2}$. It is easy to see $\bigcup\mathtt{nei}_{\emptyset}(s)\neq\emptyset$. Then, $\mathtt{nei}_{\emptyset}(s)=\{\bigcup\mathtt{nei}_{\emptyset}(s)\}$, which is a singleton. Note $\mathtt{nei}_{\mathrm{C}}(s)$ is not a singleton. Then, $\mathrm{C}\neq\emptyset$. Let $\mathrm{C}=\{a_{1},\dots,a_{n}\}$. Then, there is $\mathrm{X}_{a_{1}}\in\mathtt{nei}_{a_{1}}(s),\dots,\mathrm{X}_{a_{n}}\in\mathtt{nei}_{a_{n}}(s)$ such that $\mathrm{Y}_{1}=\mathrm{X}_{a_{1}}\cap\dots\cap\mathrm{X}_{a_{n}}$, and there is $\mathrm{X}^{\prime}_{a_{1}}\in\mathtt{nei}_{a_{1}}(s),\dots,\mathrm{X}^{\prime}_{a_{n}}\in\mathtt{nei}_{a_{n}}(s)$ such that $\mathrm{Y}_{2}=\mathrm{X}^{\prime}_{a_{1}}\cap\dots\cap\mathrm{X}^{\prime}_{a_{n}}$. Note $w\notin\mathrm{Y}_{2}$. Then $w\notin\mathrm{X}^{\prime}_{a_{i}}$ for some $a_{i}$. Note $w\in\mathrm{X}_{a_{i}}$. Then $\mathrm{X}_{a_{i}}\neq\mathrm{X}^{\prime}_{a_{i}}$. Note $u\in\mathrm{X}_{a_{i}}$ and $u\in\mathrm{X}^{\prime}_{a_{i}}$. Then $\mathrm{X}_{a_{i}}\cap\mathrm{X}^{\prime}_{a_{i}}\neq\emptyset$. We have a contradiction.
@@
By Theorem 9, clear grand-coalition-first action models can be equivalently defined in three different ways, which respectively concern all agents, all coalitions, and the grand coalition. It might be expected that the following condition, which concerns the grand coalition, is equivalent to the two conditions given in Theorem 11: for all $s\in\mathtt{ST}$, $\mathtt{nei}_{\mathtt{AG}}(s)$ is a partition of $\bigcup\mathtt{nei}_{\emptyset}(s)$. Actually, this is not the case. What follows is a counterexample.
Example 6.
Assume $\mathtt{AG}=\{a,b\}$. Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model such that:
@@
Note $\mathtt{nei}_{a}(s_{0})$ is not a partition of $\bigcup\mathtt{nei}_{\emptyset}(s_{0})$. However, it can be checked that $\mathtt{nei}_{\mathtt{AG}}(s_{0})=\{\{s_{1}\},\{s_{2}\},\{s_{3}\}\}$, which is a partition of $\bigcup\mathtt{nei}_{\emptyset}(s_{0})$.
@@
Alur, Henzinger, and Kupferman [alur_alternating-time_1998] proposed models for $\mathsf{ATL}$ based on so-called lock-step synchronous alternating transition systems. It can be shown that they are clear single-coalition-first neighborhood $\mathtt{SID}$-models.
@@
Let $\mathtt{G}\textbf{-}\mathtt{AM}$ be a clear grand-coalition-first action model. By Theorem 10, $\mathtt{G}\textbf{-}\mathtt{AM}$ is a single-coalition-first action model. It is easy to show that $\mathtt{G}\textbf{-}\mathtt{AM}$ is $z$-representable by a clear single-coalition-first neighborhood model.
@@
Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be a clear single-coalition-first neighborhood model. Define a single-coalition-first action model $\mathtt{S}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ as in the proof for Theorem 6. As shown in that proof, $\mathtt{S}\textbf{-}\mathtt{NM}$ $z$-represents $\mathtt{S}\textbf{-}\mathtt{AM}$. It is easy to check that $\mathtt{S}\textbf{-}\mathtt{AM}$ is clear. By Theorem 4, it is a grand-coalition-first action model.
@@ 6.1 Tree-like grand-coalition-first action models
Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be a grand-coalition-first action model.
@@
Specially, for every $\mathrm{C}\subseteq\mathtt{AG}$ and $\in\mathtt{ST}, ss$ is called a $\mathrm{C}$-history from $s$ to $s$.
@@
Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be a grand-coalition-first action model.
We say that $\mathtt{G}\textbf{-}\mathtt{AM}$ is a tree-like model if there is $r\in\mathtt{ST}$, called a root, such that for every $s\in\mathtt{ST}$, there is a unique $\mathtt{AG}$-history from $r$ to $s$.
Note that every tree-like model has a unique root.
@@
Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be a grand-coalition-first action model.
@@
    there is $r\in\mathtt{ST}$ such that for every $a\in\mathtt{AG}$ and $s\in\mathtt{ST}$, there is a unique $a$-history from $r$ to $s$;
@@
    there is $r\in\mathtt{ST}$ such that for every $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, there is a unique $\mathrm{C}$-history from $r$ to $s$;
@@
    there is $r\in\mathtt{ST}$ such that for every $s\in\mathtt{ST}$, there is a unique $\mathtt{AG}$-history from $r$ to $s$.
@@
Assume (1). Let $r\in\mathtt{ST}$ such that for every $a\in\mathtt{AG}$ and $s\in\mathtt{ST}$, there is a unique $a$-history from $r$ to $s$. Let $s\in\mathtt{ST}$. We want to show that there is a unique $\mathtt{AG}$-history from $r$ to $s$. Let $a\in\mathtt{AG}$.
@@
Let $\theta_{a}$ be the unique $a$-history from $r$ to $s$. Note that it is impossible that $\theta_{a}=x$ for some $x\in\mathtt{ST}$. Assume $\theta_{a}=(s_{0},\sigma_{a}^{1},s_{1},\dots,\sigma_{a}^{n},s_{n})$.
@@
Assume there is a different $\mathtt{AG}$-history $\theta^{\prime}_{\mathtt{AG}}$ from $r$ to $s$. Note that it is impossible that $\theta^{\prime}_{\mathtt{AG}}=x$ for some $x\in\mathtt{ST}$. Let $\theta^{\prime}_{\mathtt{AG}}=(t_{0},\lambda^{1}_{\mathtt{AG}},t_{1},\dots,\lambda_{\mathtt{AG}}^{m},t_{m})$.
@@
Assume (3). Let $r\in\mathtt{ST}$ such that for every $s\in\mathtt{ST}$, there is a unique $\mathtt{AG}$-history from $r$ to $s$. Let $s\in\mathtt{ST}$ and $\mathrm{C}\subseteq\mathtt{AG}$. We want to show that there is a unique $\mathrm{C}$-history from $r$ to $s$.
@@
Assume there is a $\mathrm{C}$-history $\theta_{\mathrm{C}}$ from $r$ to $r$ which is different from $r$. Note that it is impossible that $\theta_{\mathrm{C}}=x$ for some $x\in\mathtt{ST}$. Assume $\theta_{\mathrm{C}}=(s_{0},\sigma_{\mathrm{C}}^{1},s_{1},\dots,\sigma_{\mathrm{C}}^{n},s_{n})$.